Jensen Boats and Docks has 300 available slips that rent for $1,000 per season. Payments must be made in full at the start of the boating season, April 1, 2018. Slips for the next season may be reserved if paid for by December 31, 2017. Under a new policy, if payment is made by December 31, 2017, a 10% discount is allowed. The boating season ends October 31, and the marina has a December 31 year-end. To provide cash flow for major dock repairs, the marina operator is also offering a 25% discount to slip renters who pay for the 2019 season.

For the fiscal year ended December 31, 2017, all 300 slips were rented at full price. Two hundred slips were reserved and paid for the 2018 boating season, and 80 slips for the 2019 boating season were reserved and paid for.

2. How much unearned revenue would be recorded in 2017 related to the above transactions?

a. $300,000

b. $240,000

c. $540,000

d. $0

3. How much cash would have been collected in 2018 related to the above transactions?

a. $300,000

b. $240,000

c. $540,000

d. $0

4. How much earned revenue would have been recorded in 2018 related to the above transactions?

a. $240,000

b. $60,000

c.$180,000

d.$0
Answer #1

Unearned revenue refers to the money received from a customer for work that has not yet been performed.

Total cash received during fiscal year 2017:

Related to year 2017 = (300 * 1000) = $300,000

Related to year 2018 = (200 slips * $1000) * 90% = $180,000

Related to year 2019 = (80 slips * $1000) * 75% = $60,000

Ques 2. The correct answer is B

Unearned revenue to be recognized in 2017 = Amount received for years 2018 and 2019 = $ 180,000 + $ 60,000

= $ 240,000

Ques 3. The correct answer is D

No amount of cash will be received during fiscal year 2018, because the entire amount (incl. discounts) has already been paid in fiscal year 2017.

Ques 4. The correct answer is C

Earned revenue to be recorded in fiscal year 2018 = Revenue earned/related to year 2018

= $ 180,000